OpenStudy (anonymous):

Solve the polynomial equation by factoring and then using the zero product principle: 3x^4=1536x

OpenStudy (anonymous):

Subtract 1536x from both the sides, Tell me what you get??

OpenStudy (anonymous):

3x^4-1536x=0

OpenStudy (anonymous):

Yes.. take x common and tell what you got??

OpenStudy (anonymous):

factor and then i got 3x(x^3-512) which turned into x^3=512 turned into x=8 what i dont understand is how the second answer is -4+4i radical3, -4-4iradical3

OpenStudy (anonymous):

You have not use the formula...

OpenStudy (anonymous):

which fomula is that?

OpenStudy (anonymous):

See, \[(x^3 - a^3) = (x-a)(x^2 + a^2 + xa)\] After 3x(x^3 - 8^3) use this formula here and tell me what you got??

OpenStudy (anonymous):

look, this is the last question i need to turn in my online hw. if i give you a similar one can you give me the answer so i can go to sleep?

OpenStudy (anonymous):

ill learn it tomorrow but it needs to be done tonight

OpenStudy (anonymous):

brain is fried

OpenStudy (anonymous):

But you should do them on your own I am here to help how to solve and not to solve on my own...